OpenStudy (anonymous):

Write the explicit formula for the geometric sequence. a1 = -5 a2 = 20 a3 = -80

OpenStudy (anonymous):

a2/a1=??? a3/a2=???

OpenStudy (anonymous):

i dont understand how to do this?

OpenStudy (anonymous):

start by answering the 2 questions i gave you

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

so the sequence is increasing by multiplying -4 to the previous term this can be represented by \[(-4)^n\]

OpenStudy (anonymous):

what do i need to do next thought?

OpenStudy (anonymous):

though*

OpenStudy (anonymous):

\[a_n=B(-4)^{n-1}\] solve for B

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

where did you get -2 from? explicit formula \[a_n=a_1*r^{n-1}\]

OpenStudy (anonymous):

i really just dont understand at all im not getting how im supposed to figure out this problem.

OpenStudy (anonymous):

r= -4, this you determined now just substitute in all the information into the explicit formula \[a_n=a_1*r^{n-1}\]

OpenStudy (anonymous):

A geometric sequence progresses term by term by multiplying the previous term by a common ratio (that's the 'r') to get the next term. The 'n' is the number of times you multiplied by that ratio. a_1 is what you start with. If you want to find the 5th term, e.g., you multiply what you start with by the common ratio 4 times (n=5, so 4=n-1).